OpenStudy (anonymous):

There are three red marbles, four white marbles, five yellow marbles, and six black marbles in a bag. What is the probability of drawing two red marbles, if you don't replace the marble after you draw the first one? 1/51 1/6 1/36 1/54 The school play has two leading roles. The play's director decides to choose the leading actors by drawing two names from a hat that contains the names of four students who want the lead parts. You and your best friend both want to be leads. What is the probability that you will be selected first, and then your best friend will also be selected for a leading role?

OpenStudy (anonymous):

@OregonDuck

oregonduck (oregonduck):

total marbles : 3 + 4 + 5 + 6 = 18 red marble on first try = 3/18 = 1/6 red marble on second try = 2/17 now multiply......2/17 x 1/6 = 2/102 = 1/51

oregonduck (oregonduck):

so a

oregonduck (oregonduck):

1/4 * 1/4 = 1/16. The answer is D.

oregonduck (oregonduck):

:)

OpenStudy (anonymous):

:)